The coiled spring of a toy supports the weight of a child. The spring is compressed a distance of 1.6 inches by the weight of a 35-pound child. The toy will not work properly if its spring is compressed more than 6 inches. What is the weight of the heaviest child who should be allowed to use the toy? (Round your answer to two decimal places.)
Answer #1

Given,

x = 1.6 in = 0.041 m

m = 35 pound = 15.88 kg

We know that, for a vertical spring system

F(hooks) = weight

kx = mg

k = mg/x = 15.88 x 9.81/0.041 = 3780 N/m

m = kx/g

m = 3780 x 0.1524/9.8 = 59.1 kg

Hence, m = 59.1 kg = 130.3 pounds